OpenStudy (anonymous):

solve the following equation: (x - 3) /x = (x - 3) / (x - 6)

OpenStudy (anonymous):

\[\frac{x-3}{x}=\frac{x-3}{x-3}\]? unlikely

OpenStudy (anonymous):

cross multiply [x-3][x-6] and [x][x-3] expand the equations. cancel like terms and simplify youll get x=7/3

OpenStudy (anonymous):

@unknown&confused do what i said. then youll get it.

OpenStudy (anonymous):

@unknown&confused the answer is x=3, i miss typed it.

OpenStudy (anonymous):

This is basically a cross multiplication x(x-3) = (x-6)(x-6) = x^2 - 3x = X^2 - 9x + 18 the X^2's on both sides cancel and rearranging the equation will give you 18-6x = 0 Thus, x = 3 Njoy!

OpenStudy (anonymous):

no i don't think so even using the cross multiply method \[(x-3)(x-6)=x(x-3)\] \[x^2-9x+18=x^2-3\] \[-9x+18=-3\] \[-9x=-27\] \[x=3\] how do you like that, forgot that the numerator could be zero!!

OpenStudy (anonymous):

thanks everyone really appreciate the help

OpenStudy (anonymous):

Yeah! many ways to lead to the same destination! Ultimately x is 3!

OpenStudy (anonymous):